Each gallon of paint covers 400 square foot, I have to paint one side of a wall that is 8 meters tall and 40 meters long. If a food is approximately 0.3048 meters, then what is the smallest whole number of gallons I can buy and have enough paint to cover the whole wall?
Answer by stanbon(75887)

Each gallon of paint covers 400 square foot, I have to paint one side of a wall that is 8 meters tall and 40 meters long. If a foot is approximately 0.3048 meters, then what is the smallest whole number of gallons I can buy and have enough paint to cover the whole wall?
-------------
Area = 8*40 = 320 sq meters
--------------
320 m^2 (1 ft/0.3048m)^2 = 320(1 ft/0.3048 m)^2 = 3444.45 sq ft
-----
Ans:: 3444.45/400 = 8.61 gallons
----------
Cheers,
Stan H.